How many solutions does the system have?
x = -4y + 4
2x + 8y = 8

Answer by math_helper(2461)


Since the first equation can be written: x + 4y = 4
and then multiplying it by 2 gives: 2x + 8y = 8
That equation matches the 2nd one, so you really only have one equation with two unknowns.
In this case, there will be an infinite number of solutions:
2x+8y = 8
y = (1/8)(-2x + 8)
y = -(1/4)x + 1

So the solutions are all the points on the line y = -(1/4)x + 1.